OpenStudy (anonymous):

use substitution method 1/2x+1/2y=-1 1/3x-1/2y=-4 what is the solution set? i get -6,4 is this correct?

OpenStudy (anonymous):

Yes, you are correct.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

2a+4b=-24 a=20-2b i need step by step help i think I have it set up like this 2(20-2b)+4b=-24 40-4b+4b=-24 40-0b=-24 add 40 to both sides -0 divided by 16 so b has no solution

OpenStudy (anonymous):

2(20-2b) + 4b = -24 40 - 4b + 4b = -24 40 = -24 So yes, you are correct. b has no solution.

OpenStudy (anonymous):

thanks once again :)